OpenStudy (anonymous):

Change y=x^2+6x=3 from parabola equation to vertex form

OpenStudy (cwrw238):

first divide 6 ( from 6x) by 2 and write (x+3)^2 x^2 + 6x - 3 = (x + 3)^2 - a - 3 x^2 + 6x must equal (x +3)^2 - a so we need to find the value of a: (x+3)^2 = x^2 + 6x + 9 so a must be 9 x^2 + 6x - 3 = (x+3)^2 - 9 - 3 = (x+3)^2 - 12 and that s vertex form